OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. (2 points) u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

sorry again..

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

angle between two vectors say a and b \[\cos \theta = \frac{ a.b }{ |a||b| }\]

rishavraj (rishavraj):

a.b = dot product of vectors....and \[|a||b|\] means product of magnitude of vectors

OpenStudy (anonymous):

i have no idea..